# COMET: NFT Price Prediction with Wallet Profiling

### v2

@@ Abstract.
copyright: noneconference: the ACM KDD Conference; August 25–29, 2024; Barcelona, Spainccs: Information systems Data mining
@@ 2. Related Work
NFT Valuation and Price Prediction. As an emerging digital asset class, NFTs have attracted exploratory efforts that consider various influencing factors aimed at estimating their prices. We categorize them into asset valuation and price prediction, summarized in Table 1. From this table, we can observe that the asset value approaches are all for token-level NFT pricing, while the price prediction methods are for daily-level collection price prediction. The former usually estimate the current or intrinsic value of a specific NFT based on multiple relatively static factors, such as rarity (Mekacher et al., 2022), multimodal content (Costa et al., 2023), and twitter activities (Kapoor et al., 2022). However, these works lack forecasting awareness and cannot meet the requirements of investors in such a dynamic market. Existing studies on NFT price prediction mainly focus on daily collection price prediction. For example, Brunet et al. (Casale-Brunet et al., 2022) collected tweets based on keyword search, and analyzed the latent correlation between the number of tweets and the average collection price. Luo et al. (Luo et al., 2023) conducted the sentiment analysis based on tweet texts and predicted the next-day movement of collection-level price with multiple regression models. However, due to the lack of standardized problem definitions, datasets and evaluations, these approaches are hard to be fairly compared with each other. Moreover, they can only provide vague signals on which collection should be invested, while ignoring the discrepancy between a specific token and collections, e.g., the price trend of an NFT may be contrary to that of the whole collection. In addition, the influence of wallets’ multi-behaviours on the collection price is not included when modeling the predictive patterns. These shortcomings prevent the existing approaches from achieving the optima. Different from previous works, we present a practical problem definition of NFT price prediction and a wallet profiling framework to analyze the social influence on NFT prices.
@@
Graph Learning for Finance. Graph structure plays an important role in presenting financial knowledge, which facilitates understanding the correlations among financial assets (Wang et al., 2023). A notable application of this can be found in stock price prediction. Works (Li et al., 2021; Xu et al., 2021) constructed stock relationship graphs based on historical sequence correlations or stock description documents. Furthermore, studies (Li et al., 2019b; Hao et al., 2023) explored the relationship between stocks and fund managers by constructing a heterogeneous graph. Utilizing the constructed graphs, most of these studies employ graph neural networks (GNNs) to extract meaningful features for their financial tasks. Despite the prevalence of graph structure in stock market, the graph construction and its modeling method are unexplored in NFT price prediction, in which how to incorporate the multi-behaviour transactions into the graph is a challenge to be tackled.
@@
Temporal Graph Learning. As graph learning has demonstrated impressive performance on static graphs, researchers have started to explore more challenging graphs with temporal evolution. The paradigms of existing methods can be broadly categorized into several groups: timestamp embedding, parameter sequential embedding, and hidden state sequential embedding (Yang et al., 2023). Timestamp embedding-base methods (Qu et al., 2020) treat timestamps as a feature type, incorporating them alongside other features during the propagation process of GNNs. Methods based on parameter sequential embedding (Pareja et al., 2020) use RNNs to dynamically encode the parameters of GNNs across time steps. Approaches based on hidden state sequential embedding (Li et al., 2019a) encode each snapshot of data with GNNs, then employ RNNs to capture the temporal relationships between these snapshots. Although these methods are adept at capturing temporal information, they struggle with addressing challenges posed by sparse and multi-behaviour transactions in NFT domains.
@@ 3. Problem Definition
(2) $\small\hat{y}_{u}=\mathcal{F}_{u}\left(S_{u};\mathcal{F}_{c}\left(\{\mathcal{X}^{t}_{c}\mid t\in[T-H,T]\}\right)\right).$
@@ 4.1.1. Data Collection
To enhance the quality and usability of NFT-related data, we perform a series of key data preprocessing.
@@
Multimodal Content Embedding. NFT mainly consists of visual and textual contents. We employ a pretrained vision transforer (ViT) model (Dosovitskiy et al., 2021) to embed the image of NFTs, resulting in token-level visual embeddings. Regarding textual content, we concatenate various elements, including the collection-level name, description, and all token-level property keys. To derive textual embeddings, we leverage a pre-trained sentence BERT model (Reimers and Gurevych, 2019).
Token-level Rarity Calculation. As revealed in studies (Mekacher et al., 2022; He et al., 2023), the rarity of tokens strongly correlates price differences among their same collecions. For each collection $c$, we denote its propriety set as $P_{c}$. We first calculate the rarity score of each property by $r_{p}=\frac{|c|}{|p|},p\in P_{c}$, where $|c|$ denotes the number of tokens in collection $c$ and $|p_{c}|$ denotes the number of tokens that have the property $p$. Then, the rarity score of a token $i\in I_{c}$ is defined as $r_{i}=\sum_{p\in P_{c}^{i}}s_{p}$, where $P_{c}^{i}$ is the set of properties of this token.
@@ 4.2.1. Temporal Transaction Graph Construction
  • $\mathcal{R}_{t}$: There are two main types of edge, i.e., wallet to wallet and wallet to NFT, in transactions, where the former aims to capture the wallets’ social connections while the later is to model the behaviour patterns of wallets. A wallet can buy or transfer in an NFT from another wallet, which indicates two relations, i.e., sale $r_{s}$ and transfer $r_{t}$, between wallets. A wallet can mint, hold, and burn an NFT, where we denote these three relations as $r_{m}$, $r_{h}$, and $r_{b}$, respectively. Moreover, an NFT is also involved in the sale and transfer transactions between wallets, which presents four relations, i.e., sale-from, sale-to, transfer-in , transfer-out, denoted as $r_{sf}$, $r_{st}$, $r_{ti}$, and $r_{to}$, separately. For the wallet to NFT relations, we merge the NFTs to one node if they belong to the same collection. This can relieve the data sparsity problem at the NFT token-level (See Appendix A.3 for detail).
@@ 4.2.2. Global Community Detection
Within dynamic NFT ecosystem, the transaction counts of wallets exhibit a characteristic power law distribution, posing a significant challenge of data sparsity (see Appendix A.3 for detail). Notably, transfer behaviour within the NFT ecosystem tends to exhibit a close relationship between the sender and the receiver wallet (Das et al., 2022). For example, there may be different wallets registered by the same user. These wallets tend to exhibit similar market sentiments and behaviour patterns. To capture these intrinsic connections among wallets, we establish a global wallet transfer graph, based on all historical transfers within the specified time window. In this graph, nodes represent individual wallets, while edges indicate instances of transfer behaviour between these wallets. We measure the strength of the connections between wallets by considering the count of transfers as the weight of each edge. Subsequently, we employ the Louvain algorithm (Blondel et al., 2008) to conduct community detection within this graph, resulting in the identification of distinct wallet clusters. These clusters are collectively denoted as $\Phi$, and each wallet $w$ is associated with a specific cluster $\phi\in\Phi$. This approach enables us to reveal the inherent community structure among wallets, and collaboratively consider market sentiment and dynamic evolution at the community level.
@@ 4.2.3. Transaction Graph Network
(3) $\small\tilde{h}_{w,0}=\bar{h}_{w}+\mathrm{MLP}(f_{w}^{d}),\quad\forall w\in\mathcal{W},$
(4) $\small h_{c,0}=\bar{h}_{c}+\mathrm{MLP}(f_{c}^{d})+\mathrm{MLP}(f_{c}^{s}),\quad\forall c\in\mathcal{C},$
@@
Community-aware information fusion is to address the sparse transaction issue and perceive the community-driven collaboration and evolution. Specifically, we first derive the embedding $h_{\phi}$ of each cluster $\phi$ with the mean pooling method, i.e., $h_{\phi}=\frac{1}{|\phi|}\sum_{w\in\phi}\tilde{h}_{w,0},\forall\phi\in\Phi$. Then, we concatenate each wallet node embedding $h_{w}$ and its cluster embedding $h_{\phi}$ and encode through MLPs, to integrate the community-level information, obtain ${h}_{w,0}$. When the context is clear, we unifiedly use $h$ to indicate embeddings of nodes, i.e., wallets and collections, in the following presentations.
Multi-behaviour semantic aggregation is to propagate embeddings along with the constructed multi-type edges to model the wallet profile and its influence on the price of NFT collections. This process consists of multiple graph network layers, and given the $n$-th graph network layer, we consider its input node embedding as $h_{i,n},\forall i\in\mathcal{N}_{t}$. Specifically, for each distinct edge type $r\in\mathcal{R}$, we engage in adaptively semantic propagation to adjacent nodes (Veličković et al., 2018).
(5) $\small h_{i,n}^{r}=h_{i,n}+\sum_{j\in\mathcal{N}_{i}^{r}}\alpha_{i,j}h_{j,n},\quad\forall i\in\mathcal{N}_{t},\quad\forall r\in\mathcal{R}_{t},$
where $\mathcal{N}_{i}^{r}$ denotes the neigheors of node $i$ under the view of edge type $r$. $\alpha_{i,j}$ denotes adaptive weights between node $i$ and $j$. Considering there are features $h_{(i,j),n}^{r}$ for some edges, we add them with features of node $i$ and $j$ before computing the adaptive weights, i.e., $h_{(i,j),n}^{f}=h_{i,n}+h_{j,n}+h_{(i,j),n}^{r}$. For the edges without features, we set its features as zero vectors. Then, $\alpha_{i,j}$ is calculated as follows:
(6) $\small\alpha_{i,j}=\frac{\exp\left(\theta_{v,r}^{\top}\sigma(\theta_{w,r}h_{(i,j),n}^{f}\right)}{\sum_{k\in\mathcal{N}_{i}^{r}}\exp\left(\theta_{v,r}^{\top}\sigma(\theta_{w,r}h_{(i,j),n}^{f}\right)},$
@@
Here, we also introduce the residual connection to enhance the original features, while avoiding the over-smooting issues.
@@ 4.2.4. Temporal Attention Representation
After obtaining overall representations of each collection $c$ at different time steps $t$, we use a long short-term memory (LSTM) network to capture temporal patterns among sequential snapshots. We can obtain the hidden state $h^{t}_{c}$ of each collection $c$ at each timestep $t$:
(8) $\small h_{c}^{t}=\mathrm{LSTM}(h_{c}^{t-1},x_{c}^{t}),\quad\forall{t\in[T-H,T]},$
@@
(9) $\small\hat{h}_{c}=h^{T}_{c}+\!\!\!\!\!\!\sum_{t\in[T-H,T]}\!\!\!\!\!\!\alpha_{t}h_{c}^{t},\quad\alpha_{t}=\frac{\exp(\theta_{v}^{\top}\sigma(\theta_{w}(h^{t}_{c})))}{\sum_{T\in[T-H,T]}\exp(\theta_{v}^{\top}\sigma(\theta_{w}(h_{c}^{t})))}.$
@@ 4.3.2. Token-level Sale Price Predictor
For each token $u$, the features associated with each sale $s\in\mathcal{S}_{u}$ include the sale price, ETH-to-USD exchange rate, and the collection-level price on that specific day, denoted as $f_{u}^{s}$. Given the heterogeneity of token content, variations in price can exist even within the same collection. To account for these relative price relationships between individual tokens and their respective collections, we draw inspiration from the work (Sun et al., 2019) and employ a sale sequence encoder based on the Transformer (Vaswani et al., 2017) (TFEncoder). We derive the token sale embedding $h_{u}^{s}$ as follows:
(11) $\small h_{u}^{s}=\mathrm{TFEncoder}(\{f^{s}_{u}\mid s\in\mathcal{S}_{u}\}).$
@@ 5. Experiments
We evaluate COMET in tasks at both collection and token levels, comparing it to advanced baselines. We also conduct an ablation study to validate the effectiveness of each component and analyze feature importance. Furthermore, we study the impact of transaction count on performance, placed in Appendix A.5.
@@ 6. Conclusion
In this paper, we have investigated an emerging application problem, NFT price prediction. To bring clarity to this evolving field, we have provided a practical and hierarchical problem definition encompassing collection-level and token-level tasks, which cater to the diverse practical requirements of investors. we also have presented our NFT price prediction system, which includes robust data infrastructure and advanced technical solutions. Furthermore, We have proposed a novel wallet profiling framework and instantiated the COMET model, enabling us to grasp the influences on NFT prices behind transactions. Specifically, we have developed a temporal transaction graph construction method that efficiently models the multi-behavior transactions within the NFT ecosystem. Building on this, we have designed a transaction graph network, automatically extracting both temporal dynamics and structural information relevant to NFT prices. Then, we have introduced a hierarchical prediction approach to address the two-level prediction targets. Extensive experiments on two-level tasks have shown that COMET achieved the best performance, which demonstrates its effectiveness and practical applicability.

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671621isbn: 979-8-4007-0490-1/24/08ccs: Information systems Data mining
@@ 2. Related Work
NFT Valuation and Price Prediction. As an emerging digital asset class, NFTs have attracted exploratory efforts that consider various influencing factors aimed at estimating their prices. We categorize them into asset valuation and price prediction, summarized in Table 1. From this table, we can observe that the asset value approaches are all for token-level NFT pricing, while the price prediction methods are for daily-level collection price prediction. The former usually estimates the current or intrinsic value of a specific NFT based on multiple relatively static factors, such as rarity (Mekacher et al., 2022), multimodal content (Costa et al., 2023), and twitter activities (Kapoor et al., 2022). However, these works lack forecasting awareness and cannot meet the requirements of investors in such a dynamic market. Existing studies on NFT price prediction mainly focus on daily collection price prediction. For example, Brunet et al. (Casale-Brunet et al., 2022) collected tweets based on keyword search and analyzed the latent correlation between the number of tweets and the average collection price. Luo et al. (Luo et al., 2023) conducted sentiment analysis based on tweet texts and predicted the next-day movement of the collection-level price with multiple regression models. However, due to the lack of standardized problem definitions, datasets and evaluations, these approaches are hard to be fairly compared with each other. Moreover, they can only provide vague signals on which collection should be invested, while ignoring the discrepancy between a specific token and collections, e.g., the price trend of an NFT may be contrary to that of the whole collection. In addition, the influence of wallets’ multi-behaviours on the collection price is not included when modeling the predictive patterns. These shortcomings prevent the existing approaches from achieving the optima. Different from previous works, we present a practical problem definition of NFT price prediction and a wallet profiling framework to analyze the social influence on NFT prices.
@@
Graph Learning for Finance. Graph structure plays an important role in presenting financial knowledge, which facilitates understanding the correlations among financial assets (Wang et al., 2023). A notable application of this can be found in stock price prediction. Works (Li et al., 2021; Xu et al., 2021) constructed stock relationship graphs based on historical sequence correlations or stock description documents. Furthermore, studies (Li et al., 2019b; Hao et al., 2023) explored the relationship between stocks and fund managers by constructing a heterogeneous graph. Utilizing the constructed graphs, most of these studies employ graph neural networks (GNNs) to extract meaningful features for their financial tasks (Chen et al., 2024). Despite the prevalence of graph structures in stock markets, the construction and modeling methods for graphs remain unexplored in NFT price prediction. Incorporating multi-behavior transactions into graphs is a challenge that needs to be addressed.
@@
Temporal Graph Learning. As graph learning has demonstrated impressive performance on static graphs (Zheng et al., 2021; Wang et al., 2024a; Zheng et al., 2023; Ding et al., 2024), researchers have started to explore more challenging graphs with temporal evolution. The paradigms of existing methods can be broadly grouped into timestamp embedding, parameter sequential embedding, and hidden state sequential embedding (Yang et al., 2023). Timestamp embedding-base methods (Qu et al., 2020) treat timestamps as a feature type, incorporating them alongside other features during the propagation process of GNNs. Methods based on parameter sequential embedding (Pareja et al., 2020) use RNNs to dynamically encode the parameters of GNNs across time steps. Approaches based on hidden state sequential embedding (Li et al., 2019a) encode each snapshot of data with GNNs, then employ RNNs to capture the temporal relationships between these snapshots. Although these methods are adept at capturing temporal information, they struggle with addressing challenges posed by sparse and multi-behaviour transactions in NFT domains.
@@ 3. Problem Definition
(2) $\hat{y}_{u}=\mathcal{F}_{u}\left(S_{u};\mathcal{F}_{c}\left(\{\mathcal{X}^{t}_{c}\mid t\in[T-H,T]\}\right)\right).$
@@ 4.1.1. Data Collection
To improve the quality and usability of NFT-related data, we perform a series of key data preprocessing.
@@
Multimodal Content Embedding. NFT consists mainly of visual and textual content. We employ a pretrained vision transformer (ViT) model (Dosovitskiy et al., 2021) to embed the image of NFTs, resulting in token-level visual embeddings (Chen et al., 2022). Regarding textual content, we concatenate various elements, including the collection-level name, description, and all token-level property keys (He et al., 2023). To derive textual embeddings, we leverage a pre-trained sentence BERT model (Reimers and Gurevych, 2019).
Token-level Rarity Calculation. As revealed in studies (Mekacher et al., 2022; He et al., 2023), the rarity of tokens strongly correlates with price differences among their same collecions. For each collection $c$, we denote its propriety set as $P_{c}$. We first calculate the rarity score of each property by $r_{p}=\frac{|c|}{|p|},p\in P_{c}$, where $|c|$ denotes the number of tokens in collection $c$ and $|p_{c}|$ denotes the number of tokens that have the property $p$. Then, the rarity score of a token $i\in I_{c}$ is defined as $r_{i}=\sum_{p\in P_{c}^{i}}s_{p}$, where $P_{c}^{i}$ is the set of properties of this token.
@@ 4.2.1. Temporal Transaction Graph Construction
  • $\mathcal{R}_{t}$: There are two main types of edge, i.e., wallet to wallet and wallet to NFT, in transactions, where the former aims to capture the wallets’ social connections while the later is to model the behaviour patterns of wallets. A wallet can buy or transfer in an NFT from another wallet, which indicates two relations, i.e., sale $r_{s}$ and transfer $r_{t}$, between wallets. A wallet can mint, hold, and burn an NFT, where we denote these three relations as $r_{m}$, $r_{h}$, and $r_{b}$, respectively. Moreover, an NFT is also involved in the sale and transfer transactions between wallets, which presents four relations, i.e., sale-from, sale-to, transfer-in , transfer-out, denoted as $r_{sf}$, $r_{st}$, $r_{ti}$, and $r_{to}$, separately. For the wallet to NFT relations, we merge the NFTs to one node if they belong to the same collection. This can relieve the data sparsity problem at the NFT token level (See Appendix A.3 for detail).
@@ 4.2.2. Global Community Detection
Within a dynamic NFT ecosystem, the transaction counts of wallets exhibit a characteristic power law distribution, posing a significant challenge of data sparsity (see Appendix A.3 for detail). Notably, transfer behaviour within the NFT ecosystem tends to exhibit a close relationship between the sender and the receiver wallet (Das et al., 2022). For example, there may be different wallets registered by the same user. These wallets tend to exhibit similar market sentiments and behaviour patterns. To capture these intrinsic connections among wallets, we establish a global wallet transfer graph, based on all historical transfers within the specified time window. In this graph, nodes represent individual wallets, while edges indicate instances of transfer behaviour between these wallets. We measure the strength of the connections between wallets by considering the count of transfers as the weight of each edge. Subsequently, we employ the Louvain algorithm (Blondel et al., 2008) to conduct community detection within this graph, resulting in the identification of distinct wallet clusters. These clusters are collectively denoted as $\Phi$, and each wallet $w$ is associated with a specific cluster $\phi\in\Phi$. This approach enables us to reveal the inherent community structure among wallets, and collaboratively consider market sentiment and dynamic evolution at the community level.
@@ 4.2.3. Transaction Graph Network
(3) $\tilde{h}_{w,0}=\bar{h}_{w}+\mathrm{MLP}(f_{w}^{d}),\quad\forall w\in\mathcal{W},$
(4) $h_{c,0}=\bar{h}_{c}+\mathrm{MLP}(f_{c}^{d})+\mathrm{MLP}(f_{c}^{s}),\quad\forall c\in\mathcal{C},$
@@
Community-aware information fusion is to address the sparse transaction issue and perceive the community-driven collaboration and evolution. Specifically, we first derive the embedding $h_{\phi}$ of each cluster $\phi$ with the mean pooling method, i.e., $h_{\phi}=\frac{1}{|\phi|}\sum_{w\in\phi}\tilde{h}_{w,0},\forall\phi\in\Phi$. Then, we concatenate each wallet node embedding $h_{w}$ and its cluster embedding $h_{\phi}$ and encode through MLPs, to integrate the community-level information, and obtain ${h}_{w,0}$. When the context is clear, we unifiedly use $h$ to indicate embeddings of nodes, i.e., wallets and collections, in the following presentations.
Multi-behaviour semantic aggregation is to propagate embeddings along with the constructed multi-type edges to model the wallet profile and its influence on the price of NFT collections. This process consists of multiple graph network layers, and given the $n$-th graph network layer, we consider its input node embedding as $h_{i,n},\forall i\in\mathcal{N}_{t}$. Specifically, for each distinct edge type $r\in\mathcal{R}$, we engage in adaptive semantic propagation to adjacent nodes (Veličković et al., 2018).
(5) $h_{i,n}^{r}=h_{i,n}+\sum_{j\in\mathcal{N}_{i}^{r}}\alpha_{i,j}h_{j,n},\quad\forall i\in\mathcal{N}_{t},\quad\forall r\in\mathcal{R}_{t},$
where $\mathcal{N}_{i}^{r}$ denotes the neighbors of node $i$ under the view of edge type $r$. $\alpha_{i,j}$ denotes adaptive weights between node $i$ and $j$. Considering there are features $h_{(i,j),n}^{r}$ for some edges, we add them with features of node $i$ and $j$ before computing the adaptive weights, i.e., $h_{(i,j),n}^{f}=h_{i,n}+h_{j,n}+h_{(i,j),n}^{r}$. For the edges without features, we set their features as zero vectors. Then, $\alpha_{i,j}$ is calculated as follows:
(6) $\alpha_{i,j}=\frac{\exp\left(\theta_{v,r}^{\top}\sigma(\theta_{w,r}h_{(i,j),n}^{f}\right)}{\sum_{k\in\mathcal{N}_{i}^{r}}\exp\left(\theta_{v,r}^{\top}\sigma(\theta_{w,r}h_{(i,j),n}^{f}\right)},$
@@
Here, we also introduce the residual connection to enhance the original features, while avoiding the over-smoothing issues.
@@ 4.2.4. Temporal Attention Representation
After obtaining overall representations of each collection $c$ at different time steps $t$, we use a long short-term memory (LSTM) network to capture temporal patterns among sequential snapshots (Wang et al., 2021, 2024b). We can obtain the hidden state $h^{t}_{c}$ of each collection $c$ at each timestep $t$:
(8) $h_{c}^{t}=\mathrm{LSTM}(h_{c}^{t-1},x_{c}^{t}),\quad\forall{t\in[T-H,T]},$
@@
(9) $\hat{h}_{c}=h^{T}_{c}+\!\!\!\!\!\!\sum_{t\in[T-H,T]}\!\!\!\!\!\!\alpha_{t}h_{c}^{t},\quad\alpha_{t}=\frac{\exp(\theta_{v}^{\top}\sigma(\theta_{w}(h^{t}_{c})))}{\sum_{T\in[T-H,T]}\exp(\theta_{v}^{\top}\sigma(\theta_{w}(h_{c}^{t})))}.$
@@ 4.3.2. Token-level Sale Price Predictor
For each token $u$, the features associated with each sale $s\in\mathcal{S}_{u}$ include the sale price, ETH-to-USD exchange rate, and the collection-level price on that specific day, denoted $f_{u}^{s}$. Given the heterogeneity of token content, variations in price can exist even within the same collection. To account for these relative price relationships between individual tokens and their respective collections, we draw inspiration from the work (Sun et al., 2019) and employ a sale sequence encoder based on the Transformer (Vaswani et al., 2017) (TFEncoder). We derive the token sale embedding $h_{u}^{s}$ as follows:
(11) $h_{u}^{s}=\mathrm{TFEncoder}(\{f^{s}_{u}\mid s\in\mathcal{S}_{u}\}).$
@@ 5. Experiments
We evaluate COMET in tasks at both collection and token levels, comparing it to advanced baselines. We also conduct an ablation study to validate the effectiveness of each component and analyze feature importance. Furthermore, we study the impact of transaction count on performance, placed in Appendix 5.5.
@@ 5.4. Feature Importance Analysis
5.5. Effectiveness in Different Collections
To assess the performance of COMET on collections with varying transaction counts, we compare COMET with ALSTM on three representative collections (BAYC [^7] Collection address: 0xbc4ca0eda7647a8ab7c2061c2e118a18a936f13d, Doodles [^8] Collection address: 0x8a90cab2b38dba80c64b7734e58ee1db38b8992e, and VeeFrends [^9] Collection address: 0xa3aee8bce55beea1951ef834b99f3ac60d1abeeb). These collections possess similar token counts but differ significantly in transaction counts. The key distinction between COMET and ALSTM lies in COMET’s ability to harness the wallet profiling method to extract NFT price-related insights from transaction behaviors. In this analysis, we set the prediction step to 5. Figure 7 shows the improvements achieved by COMET compared to ALSTM across these collections. We observe that COMET tends to achieve more significant performance improvements in collections with higher transaction counts. This results mainly from the fact that collections with higher transaction counts typically provide richer and more diverse data for COMET. This increased data availability allows COMET to better learn and capture the underlying patterns and dynamics of the NFT market.
@@ 6. Conclusion
In this paper, we have investigated an emerging application problem, NFT price prediction. To bring clarity to this evolving field, we have provided a practical and hierarchical problem definition encompassing collection-level and token-level tasks, which cater to the diverse practical requirements of investors. we also have presented our NFT price prediction system, which includes robust data infrastructure and advanced technical solutions. Furthermore, we have proposed a novel wallet profiling framework and instantiated the COMET model, enabling us to grasp the influences on NFT prices behind transactions. Specifically, we have developed a temporal transaction graph construction method that efficiently models the multi-behavior transactions within the NFT ecosystem. Building on this, we have designed a transaction graph network, automatically extracting both temporal dynamics and structural information relevant to NFT prices. Then, we have introduced a hierarchical prediction approach to address the two-level prediction targets. Extensive experiments on two-level tasks have shown that COMET achieved the best performance, which demonstrates its effectiveness and practical applicability. Furthermore, we have deployed COMET from the perspectives of both the establishment of the data infrastructure and the integration of practical products.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant No.92370204), National Key R&D Program of China (Grant No.2023YFF0725001), Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality, and Nansha Postdoctoral Research Project.